OpenStudy (anonymous):

6[5(3−9)−1]+2 ---------------- 7(8−2)+4 The top problem is on like a fraction so its the top problem ontop of the bottom problem, but the answer isnt a fraction so i think you need to divide those answers

OpenStudy (anonymous):

Hold on, I'm gonna rewrite this using the equation button, tell me if i did it wrong

OpenStudy (anonymous):

\[\frac{ 6[5(3 - 9) - 1] + 2}{ 7(8 - 2) + 4 }\]

OpenStudy (anonymous):

Is that what it looks like?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thats what it looks like lol

OpenStudy (anonymous):

Great, now let's get to work. You know what PEMDAS is, right? Parentheses Exponents Multiplication Division Addition Subtraction

OpenStudy (anonymous):

Yes i know lol.

OpenStudy (anonymous):

@Needhelpstudying

OpenStudy (anonymous):

and i know how to do pemdas but i never get the answer right

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

First, let's do the numerator.

OpenStudy (anonymous):

the answer is 5 degrees

OpenStudy (anonymous):

i need to know how to get the answer

OpenStudy (anonymous):

\[6[5(3 - 9) - 1] + 2\]

OpenStudy (anonymous):

so parenthisis, you would start 3-9 which is 6

OpenStudy (anonymous):

The answer is 5 degrees

OpenStudy (anonymous):

To get answer, you need to use Pemdas. Break the problem up. You have to distribute first. The first thing to distribute would be 5 to (3 - 9) What is 5(3 - 9)?

OpenStudy (anonymous):

30 @Needhelpstudying

OpenStudy (anonymous):

No, 15 - 35 is -30

OpenStudy (anonymous):

But you were close

OpenStudy (anonymous):

wait what? im confused now

OpenStudy (anonymous):

5(3 - 9) right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

5 * 3 is 15 5 * -9 = -45 15 - 45 = -30

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

3 - 9 = (3) + (-9) You have to group the sign before the last term. So - 9 becomes -9.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

And then 15 - 45, remember, - 45 = -45, is -30

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

Then we get 6[-30 - 1] + 2

OpenStudy (anonymous):

then we do the -30 - 1?

OpenStudy (anonymous):

Or 6(-30 - 1) + 2

OpenStudy (anonymous):

No, then we distribute the 6 to -30 and -1

OpenStudy (anonymous):

Hold on, I might be doing this wrong. What are your answer choices?

OpenStudy (anonymous):

there are none

OpenStudy (anonymous):

i need to get the answer and type it

OpenStudy (anonymous):

@KasaCiara Alright

OpenStudy (anonymous):

So 6(-30 - 1) + 2

OpenStudy (anonymous):

Now you have to distribute. (6 * -30) (6 * -1)

OpenStudy (anonymous):

-180 and -6

OpenStudy (anonymous):

perfect what is -180 - 6?

OpenStudy (anonymous):

psst -186

OpenStudy (anonymous):

+2 = -184

OpenStudy (anonymous):

So the numerator is 184, now for the denominator

OpenStudy (anonymous):

\[7(8-2) + 4\]

OpenStudy (anonymous):

Distribute the 7 to the 8 and -2 7 * 8 = 56 7 * -2 = -24 56 - -24 = 56 + 24 = 80

OpenStudy (anonymous):

80 + 4 = 84

OpenStudy (anonymous):

denominator = 84

OpenStudy (anonymous):

numerator = -186

OpenStudy (anonymous):

\[\frac{ -186 }{ 84 }\] = -2.21428571

OpenStudy (anonymous):

Wait sorry I did the denominator wrong.

OpenStudy (anonymous):

\[7(8-2)+4\] 7 * 8 = 56. 7 * -2 = -14 56 - 14 = 42 42 + 4 = 46

OpenStudy (anonymous):

-186 / 46 = -4.04347826 Which can be rounded down to 4

OpenStudy (anonymous):

That is the correct answer

OpenStudy (anonymous):

@KasaCiara